# Robust Quantum Control via Multipath Interference for Thousandfold Phase Amplification in a Resonant Atom Interferometer

### v1

@@ Abstract
Quantum sensing is a powerful tool for fundamental physics tests [1, 2], which include precision measurements of oscillating signals [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Resonant [21, 22, 23, 24, 25, 26] light-pulse atom interferometers [27, 28, 29, 30, 31, 32, 19, 33, 34, 20, 35], a class of light-pulse atom interferometers [36, 37, 38, 39], employ sequential mirror pulses to induce periodic reversals in the directions of the interferometer arms, intertwining their spacetime trajectories into loops repeating at a tunable target frequency (Fig. 1). Resonant atom interferometers have been proposed as powerful sensors of various oscillatory signals, including those from gravitational waves [35, 40, 14, 41, 42], dark matter [7, 8, 9, 14, 15, 34], dark energy [19], accelerations [8, 20, 29], and rotations [20]. Increasing the number of loops enhances interferometer sensitivity at the target frequency and suppresses off-resonant noise, akin to applying lock-in detection [43, 44, 45, 46, 47, 48, 49, 50, 51] and dynamical decoupling [52, 53, 54, 55, 56, 57] to matter waves.
@@
A critical challenge impeding the full metrological potential of resonant light-pulse atom interferometers is the adverse impact of mirror pulse infidelities, which limit the total number of loops that can be performed [20, 29, 33]. With imperfect pulse efficiency, each atom’s wavefunction diverges into many ‘stray’ paths that deviate from the central two interferometer arms (Fig. 1), thereby diminishing the population contributing to the signal and/or introducing deleterious effects through multipath interference [58, 32] [^1] Note that in certain scenarios, multipath interference [117, 118, 119] has been used to destructively reduce losses in other types of atom interferometers [120, 90].
@@
We instead introduce the method of making the entire interferometer sequence robust to imperfections in individual mirror operations by leveraging QOC [91, 92, 93, 94, 95, 96, 97, 98, 99, 100] in a novel way to advantageously harness multipath interference between numerous stray trajectories. As a demonstration, we utilize this method to enhance the robustness of a resonant single-photon clock atom interferometer to pulse infidelities, achieving 50-fold improvement in the resonant signal amplification—for a total of more than 500 loops—in spite of an individual pulse transfer efficiency of $\approx 90\%$, representative of what one might encounter for a long-baseline atom interferometer [15] such as MAGIS-100 [13]. Moreover, we elucidate how the interplay between spontaneous emission and many-pulse sequences can generate spurious interference signals with surprisingly high visibility that obscure the signals of interest, and demonstrate an optimization approach to circumvent this effect. These findings may be valuable for improving the performance of a variety of quantum sensors which are vulnerable to the effects of spontaneous emission [101, 102]. Our general approach is not constrained by the specific type of atom optics, and we anticipate that it will enable higher levels of precision across various atom interferometer designs, including those based on multi-photon transitions.
Apparatus.– A horizontal laser beam [103] tuned to the 689 nm ${}^{1}S_{0}-^{3}P_{1}(m_{j}=0)$ intercombination line performs interferometry on a cloud of 88Sr atoms released from a 3D magneto-optical trap. The intermediate 21.6 $\mu$s lifetime of this transition is long enough to be suitable for clock atom interferometry and enables high (MHz-scale) Rabi frequencies with Watt-scale interferometer beams [67, 68], which is advantageous for measuring up to MHz-scale signals, such as in dark matter searches [3]. After an interferometer sequence, a 461 nm push beam spatially separates the ground state from the excited state; then the two populations are read out via fluorescence detection [67, 68]. The interferometer fringe visibility ($v$) and phase ($\Delta\phi$) are extracted by fitting the normalized excited state population ($P_{e}$) as a function of the final beamsplitter phase ($\phi_{b}$), to the expression $\frac{1}{2}(1+v\cos[\Delta\phi+\phi_{b}])$. An M-Labs ARTIQ timing system controls the power, phase, and duration of each pulse via an acousto-optic modulator. The ratio of the laser beam waist to the atom cloud size ($\approx 3$) and the ratio of the Doppler detuning spread to Rabi frequency ($\approx 0.1$) used in this work [104] are representative of the ratios anticipated for MAGIS-100.
@@
Figure 2: Introduction to the N=8-pulse sequence with optimized phases $(\frac{3\pi}{8},\frac{7\pi}{8},\frac{15\pi}{8},\frac{11\pi}{8},\frac{11\pi}{8},\frac{15\pi}{8},\frac{7\pi}{8},\frac{3\pi}{8})$. A comparison of the spatial spread of trajectories, the laser phases $\phi_{L}$, and experimentally collected interferometer fringes in 16-loop unoptimized (a) and optimized (b) sequences. (c) Interferometer fringe visibility as a function of loop number for optimized (black) and unoptimized (blue) sequences. The red curve indicates the spontaneous emission limit. Error bars are generated via standard error from fits of the interferometer fringes. (d) The phase amplification of the optimized sequence observed after 504 loops. Each data point represents an average derived from five fringe scans, with error bars calculated from the standard deviation. (e) Leveraging the enhanced interferometer phase sensitivity to resolve the bit discretization of the phase control system in a 504 loop interferometer sequence. Red horizontal lines indicate expected phase shifts due to discretization effects. Average values and error bars, generated from standard error of the mean, are evaluated from nine fringe scans. (f) Data of the self-normalized resonant response vs. offset ($\delta f$) from resonant frequency ($f_{R}$) with theory curves for the optimized sequence for loop numbers 8 (red), 64 (blue), and 256 (yellow). Error bars are generated via standard error from fits of the interferometer fringes.
@@
Open-loop optimization with QOC. We use a semi-classical model [105, 106, 107] to compute the spread of trajectories in a multi-loop interferometer under imperfect mirror operations for a particular choice of laser phases [104]. We apply QOC to these interferometers by selecting the laser phases of the mirror pulses as control parameters [82, 53] and defining a cost function aimed at minimizing the spatial spread of trajectories throughout the interferometer cycle [104]. Our approach leverages the fact that stray trajectories spawned by one $\pi$-pulse can spatially overlap and interfere with those spawned by previous $\pi$-pulses. The phase of the laser is imprinted on kicked trajectories, which we find can facilitate destructive interference among paths diverging from the central interferometer arms and constructive interference among stray paths that rejoin them, as visualized in Fig 2 (a-b).
@@
We define an $N$-dimensional optimization problem where the phases of the first $N$ mirror pulses repeat $L/N$ times to compose an $L$-loop interferometer. We arrive at the sequence characterized in Figures 2 and 3 by tuning the phases of the mirror pulses to minimize the spread of trajectories over a set of $N=8$ laser phases repeated 8 times. Our optimization approach not only yields sequences that prove effective experimentally, but also provides a visual explanation of why certain sequences outperform others. For example, we notice that for the case of $L=8$, the universally robust (UR) UR-8 dynamical decoupling sequence, originally derived under an entirely different framework [57], provides near-optimal results per our cost function. Following the metrics used in [57] would suggest that for $L$ loops, it is best to implement the corresponding UR-$L$ sequence. For larger $L$, however, we observe that applying UR-$L$ to a resonant atom interferometer leads to a reduced interferometer phase sensitivity to resonantly oscillating signals. The semi-classical model visually illustrates the reason for this: The further a trajectory strays from the central arms, the less phase information it carries about the resonantly oscillating signal due to the fewer imprinted laser phases. The UR-$L$ sequence leads to a spread of spatial trajectories which increases with $L$, thereby reducing overall interferometer sensitivity to oscillating signals as $L$ increases. Applying the spatial spread cost optimization for $N=8$ at large $L$ produces pulse phases which happen to coincide with a repeated UR-8 pattern—outperforming UR-$L$ in our application—and optimizing over $N=16$ phases produces sequences distinct from the UR-$N$ formula owing to the reduced phase sensitivity associated with the UR-$N$ sequences at high $N$. Extending our optimization from $N=8$ to $N=16$ enhances the interferometer visibility by 1.5x in a 496-loop interferometer, while maintaining sensitivity to oscillating signals [104]. The cost function can be tuned to adjust how strongly the population in stray trajectories is penalized as a function of distance from the central arms. Limiting the spread of trajectories could be useful for accommodating constraints on the interferometer size, like in cases where the interferometer needs to fit inside a satellite [35] or a magnetically shielded region [9].
@@
We demonstrate that the phase response of the interferometer with the optimized sequence and imperfect mirror pulses is comparably sensitive to resonantly oscillating signals as the expected response of an interferometer with perfect atom-optics pulses at $L=504$ loops. We apply a signal in the form of an oscillating laser phase, $\delta\phi\sin(\omega t_{i})$, added to $i$th mirror pulse of the optimized sequence, where the pulse occurs at time $t_{i}$. An analogous signal could emerge from motion of the atom relative to the laser beam, which results in a change in phase associated with each atom-light interaction [104]. Moreover, when transforming into a frame that rotates with the oscillating laser phase, the dynamics of the system are equivalent to one in which there are oscillations in the atom’s resonance frequency [64], e.g. due to dark matter [7]. In Fig. 2(d), the interferometer phase response to a resonantly oscillating signal is in close agreement with the expected signal amplification factor of $2L\approx 1000$, indicated by the red line. In Fig. 2(e), we leverage this thousandfold amplification to resolve the digitization inherent to the AD9910 Direct Digital Synthesizer used to modulate the interferometer beam’s phase. The 16 bit phase resolution introduces observable discrete jumps of $\frac{2\pi}{2^{16}}=96\;\mu$rad, attributed to the toggling of the least significant bit. Fig. 2(f) shows the narrowing of the resonant response as a function of signal frequency as loop number is increased, consistent with the theoretical response function of an interferometer with perfect atom-optics operations [104].
Figure 3: Susceptibility of the $N=8$ optimized sequence to detuning (a) and Rabi frequency error (b) at 504 loops, as determined by experiment (solid dots), and simulation (shaded bands, width indicating uncertainty). The interferometer visibilities (blue) are normalized to their maximum value, and the interferometer phase is displayed in red. Each data point is the average obtained from 5 individual fringe scans with the error bar being the standard deviation. The detuning is plotted relative to the estimated Rabi frequency for atoms at the center of the cloud [104].
@@
Figure 4: (a) Simplified depiction of a Bloch sphere from a top-down view, composed of only atoms that have undergone spontaneous emission decay. The black arrow indicates the axis of state rotation based on the pulse phase, the yellow represents the semicircular arc formed due to atom decay at various times during the corresponding pulse and subsequent state rotation during the remainder of the pulse, and the blue represents arcs from previous pulses. (i,ii,iii) Bloch sphere after the first, second, and third mirror pulses, respectively, in the alternating ±$\pm\pi/2$ phase sequence, showing a non-zero phase formed cumulatively from decayed atoms. (iv,v,vi) Bloch sphere after the first, forth, and eighth mirror pulses, respectively, in the optimized phase sequence (see Fig. 2), which achieves a symmetric distribution of decayed atoms (arrow removed for visual clarity in vi)). (b) Simulation results of fitness values for pulse sequences over the $\phi_{1}$-$\phi_{2}$ parameter space (see main text). The inset details an example optimization trajectory for the closed-loop search method (“$\times$”: starting point, “$+$”: open-loop optimization result). (c) Visibility and phase shift for the alternating ±$\pm\pi/2$ phase sequence over 128 loops vs. amplitude $\delta\phi$ of a resonantly oscillating signal. For $\left|\delta\phi\right|\lesssim 0.1$, visibility markedly increases, but the phase shift shows a largely insensitive response. For example, when shifting from $\delta\phi=0$ to the neighboring data point with $\delta\phi=\pi/256$, the phase shift for a fully phase sensitive sequence–such as our optimized sequence–would change by approximately $\pi$. Here, the phase shift only changes by $\mathcal{O}(100\;\text{mrad})$. Data are consistent with simulations that include spontaneous emission decay (solid curves) and diverge from those excluding it (dashed curves).
In Fig. 3, we examine the resilience of the optimized sequence to deviations in overall laser detuning and Rabi frequency by deliberately introducing errors in these parameters on the order of 10% of the nominal Rabi frequency. Despite these variations, the interferometer phase remains robust within approximately ±$\pm 100$ mrad and the visibility remains above 50% of its peak value.
@@
Spurious interference from spontaneous emission. Spontaneous emission occurs stochastically throughout the interferometer sequence [108], causing atoms to lose all prior phase information stored between the ground and excited states upon decay. Spontaneous emission is commonly treated as contributing to an incoherent background in atom interferometry [109]. However, atoms undergoing spontaneous emission during a pulse can re-emerge into a coherent superposition due to the remaining pulse duration, with the phase of the laser imprinted onto the phase of the transferred population. We find that certain repeating phase sequences can synchronize the re-emerging atomic coherence from one pulse with that from previous pulses, leading to the build-up of an interference signal with a well-defined cumulative phase that is largely insensitive to oscillating signals of interest. At high loop numbers, this spurious interference pattern sometimes becomes dominant, obscuring the intended signal (Fig.4). Furthermore, we find that imposing specific constraints on the pulse sequence can suppress this spurious interference by making the final phase distribution of the decayed atoms average to zero–for example, groups of $N=8$ mirror pulses with the pattern $\phi_{1},\phi_{2},\phi_{2}+\pi,\phi_{1}+\pi,\phi_{1}+\pi,\phi_{2}+\pi,\phi_{2},\phi_{1}$ for arbitrary phases $\phi_{1}$ and $\phi_{2}$. We note that the open-loop optimized sequence obeys this constraint. Fig. 4 depicts both a phase sequence that leads to cumulative phase buildup involving mirror pulses with phase alternating between ±$\pm\pi/2$ [110], and another where we have successfully mitigated this effect using the above constraint. To conceptually illustrate cumulative phase build-up, we consider the interferometer in momentum space so that we can interpret our measurement as an average over atoms with a distribution of momenta $p$, with each atom acting as a two-level system corresponding to the coupling of states |$\ket{{}^{1}S_{0},p}$ and |$\ket{{}^{3}P_{1},p+\hbar k}$. This allows a Bloch sphere [64] visualization and optical Bloch equation (OBE) simulations [104].
Closed-loop optimization. In order to address experimental errors beyond effects captured by simulations, we survey the $\phi_{1}$-$\phi_{2}$ parameter space under the constraint described above by implementing closed-loop optimization, where optimization steps are determined by real-time experimental data [111, 112, 113, 114, 115]. We maximize a fitness function defined as $f=v\cos^{4}\left(\Delta\phi_{0}/2\right)$, where $v$ is the measured visibility and $\Delta\phi_{0}$ is the measured phase shift without an applied signal [104]. We begin the optimization near a point that our OBE simulations suggests maximizes the fitness function. The optimization trajectory converges to a pulse sequence consistent with simulation results and with the results of the open-loop optimization (Fig. 4(b)).
@@
Outlook. In future work, we aim to extend our optimization approach to interferometers using various types of atom optics and geometries, including those employing broadband dynamical decoupling [116] or multiloop configurations with increased momentum splittings, which in some cases can enhance sensitivity [35]. We have already experimentally demonstrated the adaptability of our open-loop optimization to improve robustness in the latter case, which will be discussed in a future manuscript.

### v2

@@ Abstract
Quantum sensing is a powerful tool for both fundamental physics and practical measurements [1, 2, 3], including the detection of oscillating signals [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Resonant [22, 23, 24, 25, 26, 27] light-pulse atom interferometers [28, 29, 30, 31, 32, 33, 20, 34, 35, 21, 36], a class of light-pulse atom interferometers [37, 38, 39, 40], employ sequential mirror pulses to periodically reverse the directions of the interferometer arms, creating loops that repeat at a tunable target frequency (see Fig. 1). These resonant (or multiloop) atom interferometers can be used as highly sensitive detectors for various oscillatory signals [36, 41, 15, 42, 43, 8, 9, 10, 15, 16, 35, 20, 9, 21, 30].
Oscillatory signals, whether intrinsic or engineered through experimental design, span a wide range of applications within atom interferometry, making resonant atom interferometers highly versatile for many of the field’s current research objectives. The target frequency at which the interferometer is maximally sensitive is easily tunable by adjusting the temporal spacing between mirror pulses, with the upper frequency limit constrained by the pulse repetition speed, which is in turn determined by the available Rabi frequency. Increasing the number of loops enhances sensitivity at the target frequency while suppressing off-resonant noise [36, 28, 14], akin to applying lock-in detection [44, 45, 46, 47, 48, 49, 50, 51, 52] and dynamical decoupling [53, 54, 55, 56, 57, 58] to matter waves.
Multiloop atom interferometers are powerful tools for probing fundamental physics. One major application is gravitational wave detection in the frequency range of 0.03 - 3 Hz, which complements the sensitivity of laser interferometers like LIGO and LISA [15, 16]. Large-scale atom interferometric detectors are being developed worldwide by an interdisciplinary community [15, 16, 14, 18, 59, 60, 61, 19, 17]. Resonant atom interferometry has been proposed to optimize the sensitivity of these interferometers to gravitational waves [36, 41, 15, 42, 43]. Another application is searches for wavelike dark matter candidates, which are considered among the most promising dark matter models [4]. Signatures of these candidates would manifest as an oscillatory signal in an atom interferometer, and resonant atom interferometers are expected to significantly enhance sensitivity across a broad frequency range from 10 mHz to several MHz [8, 9, 10, 15, 16, 35, 30].
Resonant atom interferometers also have practical applications. These interferometers could be used in geophysical studies to measure oscillating seismically-induced accelerations [21], either from direct vibrations of the Earth or from density fluctuations caused by seismic waves [19]. The latter measurements are not only of intrinsic geophysical interest [19], but are also crucial for characterizing noise backgrounds in both laser- and atom-based gravitational wave detectors [19, 62, 63, 64]. Multiloop interferometers are also useful in inertial navigation because they act as selective sensors for rotations or accelerations [28, 33] and have the potential to amplify signals in atomic gyroscopes [34, 29].
In some cases, experiments can be designed to cause otherwise DC signals to manifest as oscillatory in order to enable the signal amplification and noise suppression benefits of resonant detection. One example is an ongoing dark energy search experiment [20], which employs a spatially modulated source mass to induce a modulation in the dark energy field in the frame of the atoms. A similar approach has been suggested for measuring the Newtonian gravitational constant [21]. The improvement to resonant atom interferometry presented in our work may encourage the further development of experiments which render DC signals oscillatory.
@@
A critical challenge impeding the full metrological potential of resonant interferometers is the adverse impact of mirror pulse infidelities, which limit the total number of loops that can be performed [21, 30, 34]. With imperfect pulse efficiency, each atom’s wavefunction diverges into many ‘stray’ paths that deviate from the central two interferometer arms (Fig. 1), thereby diminishing the population contributing to the signal and/or introducing deleterious effects through multipath interference [65, 33] [^1] Note that in certain scenarios, multipath interference [125, 126, 127] has been used to destructively reduce losses in other types of atom interferometers [128, 94].
@@
We instead introduce the method of making the entire interferometer sequence robust to imperfections in individual mirror operations by leveraging QOC [95, 96, 97, 98, 99, 100, 101, 102, 103, 104] in a novel way to advantageously harness multipath interference between numerous stray trajectories. As a demonstration, we utilize this method to enhance the robustness of a resonant single-photon clock atom interferometer (using the 689 nm ${}^{1}S_{0}-^{3}P_{1}(m_{j}=0)$ intercombination line of 88Sr) to pulse infidelities, achieving 50-fold improvement in the resonant signal amplification—for a total of more than 500 loops—in spite of an individual pulse transfer efficiency of $\approx 90\%$, representative of what one might encounter for a long-baseline atom interferometer [16] such as MAGIS-100 [14]. Our general approach is not constrained by the specific type of atom optics, and we anticipate that it will enable higher levels of precision across various atom interferometer designs, including those based on multi-photon transitions.
Moreover, we elucidate how the interplay between spontaneous emission and many-pulse sequences can produce spurious interference signals with unexpectedly high visibility, which can obscure the signals of interest. To address this, we present an optimization approach that effectively mitigates this effect. These findings could significantly enhance the performance of quantum sensors susceptible to spontaneous emission, especially as advancements in quantum control techniques enable increasingly complex pulse sequences [105, 106]. For instance, many-pulse sequences can improve the performance of atomic clocks used in gravitational wave and dark matter detection [107, 108], as well as optical spectrum analysis [109], and the effect we have identified will likely be crucial to consider in future applications of atomic clocks in these areas. Notably, we observe a substantial spurious interference signal using the same pulse phase sequence reported for the optical spectrum analyzer in [109], suggesting that similar effects may arise when extending this work to larger pulse numbers.
@@
Figure 2: Introduction to the N=8-pulse sequence with optimized phases $(\frac{3\pi}{8},\frac{7\pi}{8},\frac{15\pi}{8},\frac{11\pi}{8},\frac{11\pi}{8},\frac{15\pi}{8},\frac{7\pi}{8},\frac{3\pi}{8})$. A comparison of the spatial spread of trajectories, the laser phases $\phi_{L}$, and experimentally collected interferometer fringes (excited state population $P_{e}$ vs. final beam splitter phase $\phi_{B}$) in 16-loop unoptimized (a) and optimized (b) sequences. (c) Interferometer fringe visibility as a function of loop number for optimized (black) and unoptimized (light blue) sequences. The solid red curve indicates the spontaneous emission limit. Error bars are generated via standard error from fits of the interferometer fringes. (d) The phase amplification of the optimized sequence observed after 504 loops. Each data point represents an average derived from five fringe scans, with error bars calculated from the standard deviation. Inset shows the ability of the enhanced phase sensitivity to resolve bit discretization in the phase control for 504 loops. Gray lines indicate expected phase shifts for the minimum laser phase modulation, limited by bit discretization. Average values and error bars, generated from standard error of the mean, are evaluated from 45 fringe scans. (e) Data of the self-normalized resonant response vs. offset ($\delta f$) from resonant frequency ($f_{R}$) with theory curves for the optimized sequence for loop numbers 8 (dashed red), 64 (dotted blue), and 256 (solid yellow). Error bars are generated via standard error from fits of the interferometer fringes. (f) The science chamber with the orientations of the atom beam, MOT beams, the interferometer beam, and the axis along which the imaging camera points (the ‘Imaging Axis’). A third MOT beam, along with push and fluorescence beams, come in and out of the page. The atom cloud is prepared in the center of the chamber, and the interferometer beam is perpendicular to the direction of gravity.
@@
Open-loop optimization with QOC. We use a semi-classical model [110, 111, 112] to compute the spread of trajectories in a multiloop interferometer under imperfect mirror operations for a particular choice of laser phases [113]. We apply QOC to these interferometers by selecting the laser phases of the mirror pulses as control parameters [86, 54] and defining a cost function aimed at minimizing the spatial spread of trajectories throughout the interferometer cycle [113]. Our approach leverages the fact that stray trajectories spawned by one $\pi$-pulse can spatially overlap and interfere with those spawned by previous $\pi$-pulses. The phase of the laser is imprinted on kicked trajectories, which we find can facilitate destructive interference among paths diverging from the central interferometer arms and constructive interference among stray paths that rejoin them, as visualized in Fig 2 (a-b).
@@
We define an $N$-dimensional optimization problem where the phases of the first $N$ mirror pulses repeat $L/N$ times to compose an $L$-loop interferometer. We arrive at the sequence characterized in Figures 2 and 5 by tuning the phases of the mirror pulses to minimize the spread of trajectories over a set of $N=8$ laser phases repeated 8 times (see End Matter for further details). Our optimization approach not only yields sequences that prove effective experimentally, but also provides a visual explanation of why certain sequences outperform others. For example, we notice that for the case of $L=8$, the universally robust (UR) UR-8 dynamical decoupling sequence, originally derived under an entirely different framework [58], provides near-optimal results per our cost function. Following the metrics used in [58] would suggest that for $L$ loops, it is best to implement the corresponding UR-$L$ sequence. For larger $L$, however, we observe that applying UR-$L$ to a resonant atom interferometer leads to a reduced interferometer phase sensitivity to resonantly oscillating signals. The semi-classical model visually illustrates the reason for this: The further a trajectory strays from the central arms, the less phase information it carries about the resonantly oscillating signal due to the fewer imprinted laser phases. The UR-$L$ sequence leads to a spread of spatial trajectories which increases with $L$, thereby reducing overall interferometer sensitivity to oscillating signals as $L$ increases (see Fig. 4 in End Matter). Applying the spatial spread cost optimization for $N=8$ at large $L$ produces pulse phases which happen to coincide with a repeated UR-8 pattern—outperforming UR-$L$ in our application—and optimizing over $N=16$ phases produces sequences distinct from the UR-$N$ formula owing to the reduced phase sensitivity associated with the UR-$N$ sequences at high $N$. Extending our optimization from $N=8$ to $N=16$ enhances the interferometer visibility by 1.5x in a 496-loop interferometer, while maintaining sensitivity to oscillating signals [113].
@@
We demonstrate that the phase response of the interferometer with the optimized sequence and imperfect mirror pulses is comparably sensitive to resonantly oscillating signals as the expected response of an interferometer with perfect atom-optics pulses at $L=504$ loops. We apply a signal in the form of an oscillating laser phase, $\delta\phi\sin(\omega t_{i})$, added to $i$th mirror pulse of the optimized sequence, where the pulse occurs at time $t_{i}$. An analogous signal could emerge from motion of the atom relative to the laser beam, which results in a change in phase associated with each atom-light interaction [113]. Moreover, when transforming into a frame that rotates with the oscillating laser phase, the dynamics of the system are equivalent to one in which there are oscillations in the atom’s resonance frequency [71], e.g. due to dark matter [8]. In Fig. 2(d), the interferometer phase response to a resonantly oscillating signal is in close agreement with the expected thousandfold phase amplification factor, indicated by the black line. We leverage this $2L\approx 1000$ phase amplification to resolve the digitization inherent to the AD9910 Direct Digital Synthesizer used to modulate the interferometer beam’s phase. The 16 bit phase resolution yields a minimum modulation amplitude of $\frac{2\pi}{2^{16}}=96\;\mu$rad, corresponding to toggling the least significant bit. The inset of panel (d) shows that the error in the measured interferometer phase is smaller than the amplified phase shift from bit flips in a 504-loop sequence, allowing the phase shift from bit flips to be clearly resolved. Fig. 2(e) shows the narrowing of the resonant response as a function of signal frequency as loop number is increased, consistent with the theoretical response function of an interferometer with perfect atom-optics operations [113]. Fig. 2(f) illustrates the experimental apparatus (see End Matter for further apparatus details). The resilience of the optimized sequence to deviations in overall laser detuning and Rabi frequency are discussed in the End Matter.
@@
Figure 3: (a) Simplified depiction of a Bloch sphere from a top-down view, composed of only atoms that have undergone spontaneous emission decay. The black arrow indicates the axis of state rotation based on the pulse phase, the yellow (light) represents the semicircular arc formed due to atom decay at various times during the corresponding pulse and subsequent state rotation during the remainder of the pulse, and the blue (dark) represents arcs from previous pulses. (i,ii,iii) Bloch sphere after the first, second, and third mirror pulses, respectively, in the alternating ±$\pm\pi/2$ phase sequence, showing a non-zero phase formed cumulatively from decayed atoms. (iv,v,vi) Bloch sphere after the first, forth, and eighth mirror pulses, respectively, in the optimized phase sequence (see Fig. 2), which achieves a symmetric distribution of decayed atoms (arrow removed for visual clarity in vi)). (b) Simulation results of fitness values for pulse sequences over the $\phi_{1}$-$\phi_{2}$ parameter space. The inset details an example optimization trajectory for a closed-loop search method (“$\times$”: starting point, “$+$”: open-loop optimization result) (see End Matter for further details). (c) Visibility and phase shift for the alternating ±$\pm\pi/2$ phase sequence over 128 loops vs. amplitude $\delta\phi$ of a resonantly oscillating signal. For $\left|\delta\phi\right|\lesssim 0.1$, visibility markedly increases, but the phase shift shows a largely insensitive response. For example, when shifting from $\delta\phi=0$ to the neighboring data point with $\delta\phi=\pi/256$, the phase shift for a fully phase sensitive sequence–such as our optimized sequence–would change by approximately $\pi$. Here, the phase shift only changes by $\mathcal{O}(100\;\text{mrad})$. Data are consistent with simulations that include spontaneous emission decay (solid curves) and diverge from those excluding it (dashed curves).
@@
Spurious interference from spontaneous emission. Spontaneous emission occurs stochastically throughout the interferometer sequence [114], causing atoms to lose all prior phase information stored between the ground and excited states upon decay. Spontaneous emission is commonly treated as contributing to an incoherent background in atom interferometry [115], though it has been noted that spontaneous emission does not always destroy atom interference [116, 117]. However, atoms undergoing spontaneous emission during a pulse can re-emerge into a coherent superposition due to the remaining pulse duration, with the phase of the laser imprinted onto the phase of the transferred population. We find that certain repeating phase sequences can synchronize the re-emerging atomic coherence from one pulse with that from previous pulses, leading to the build-up of an interference signal with a well-defined cumulative phase that is largely insensitive to oscillating signals of interest. At high loop numbers, this spurious interference pattern sometimes becomes dominant, obscuring the intended signal (Fig. 3). Furthermore, we find that imposing specific constraints on the pulse sequence can suppress this spurious interference by making the final phase distribution of the decayed atoms average to zero–for example, groups of $N=8$ mirror pulses with the pattern $\phi_{1},\phi_{2},\phi_{2}+\pi,\phi_{1}+\pi,\phi_{1}+\pi,\phi_{2}+\pi,\phi_{2},\phi_{1}$ for arbitrary phases $\phi_{1}$ and $\phi_{2}$. We note that the open-loop optimized sequence obeys this constraint. Fig. 3 depicts both a phase sequence that leads to cumulative phase buildup involving mirror pulses with phase alternating between ±$\pm\pi/2$ [109], and another where we have successfully mitigated this effect using the above constraint. To conceptually illustrate cumulative phase build-up, we consider the interferometer in momentum space so that we can interpret our measurement as an average over atoms with a distribution of momenta $p$, with each atom acting as a two-level system corresponding to the coupling of states |$\ket{{}^{1}S_{0},p}$ and |$\ket{{}^{3}P_{1},p+\hbar k}$. This allows a Bloch sphere [71] visualization and optical Bloch equation (OBE) simulations [113].
@@
Outlook. Our general optimization framework is not specific to any particular interferometer geometry or atom optics type, thus opening the new avenue of adapting it to enhance various classes of atom interferometers, and matter-wave interferometers more generally. In future research, we plan to apply our optimization method to interferometers utilizing different atom optics and geometries, including those employing broadband dynamical decoupling [108]. We also aim to explore both single and multiloop interferometer configurations with larger momentum splittings, where the additional momentum enhances sensitivity [36]. We have already experimentally observed the adaptability of our open-loop optimization to improve robustness in the latter case, which will be discussed in a future manuscript.